In the following,

\sum\limits_{m=1}^{135} 3m∑ m = 1 135 3 m

what is the value of first term?
Answer #1

13 Ś 3m = 361) + 3(2) +3(3)+--- +3(035) = 3 + 6 + 9 + -- it 395 the first term of summation we know that the expansion of the13 Ś 3m = 361) + 3(2) +3(3)+--- +3(035) = 3 + 6 + 9 + -- it 395 the first term of summation we know that the expansion of the